OpenStudy (anonymous):

What is the solution of the matrix equation? [9,4_2,1]X=[-9,-6_-1,-8]

OpenStudy (anonymous):

is it a 4X2 matrix? can you use the draw, and trying to write it clearly so it would be easy to understand your problem

OpenStudy (anonymous):

I don't know how to use draw helpme1.2

OpenStudy (nurali):

\[\left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]X=\left[\begin{matrix}-9 & -6 \\ -1 & -8\end{matrix}\right]\]

OpenStudy (nurali):

Ur Question is this?

OpenStudy (anonymous):

YES! Thank you nurali

OpenStudy (nurali):

\[X = BA ^{-1}\]

OpenStudy (anonymous):

Ok can you walk me through on how to get the answer? Nurali

OpenStudy (nurali):

\[A=\left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]\]

OpenStudy (nurali):

\[A ^{-1}=\frac{ \left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right] }{ 9-8 }\]

OpenStudy (nurali):

\[A ^{-1}=\left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right]\]

OpenStudy (nurali):

\[X=\left[\begin{matrix}-9 & -6 \\ -1 & -8\end{matrix}\right]*\left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right]\]

OpenStudy (anonymous):

Thank you so much!!!!!! Nurali!

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

would it be [9,26_9,1]?

OpenStudy (nurali):

No

OpenStudy (anonymous):

hm...

OpenStudy (nurali):

\[X=\left[\begin{matrix}3 &-18 \\ 15 & -68\end{matrix}\right]\]

OpenStudy (anonymous):

thats not in the choices...